# The trimer paradox: the effect of stiff constraints on equilibrium distributions in overdamped dynamics

### v1

@@ III.1 Computation of the shape term
Since the eigenvectors of $\bm{\mathsf{B}}$ are orthogonal and zero eigenvectors lie inside the tangent space $T\mathcal{M}$, we can compute $|\bm{\mathsf{H}}|$ from the product of the non-zero eigenvalues of $\bm{\mathsf{B}}$. We can use the knowledge of the tangent space of $\mathcal{M}$ to help us compute the eigenvalues of $\bm{\mathsf{B}}$. If our confining function is of the form $W(\bm{q})=\sum_{i}P_{i}^{2}$, we can find vectors that lie in the normal space $\mathcal{M}^{\perp}$ [15, proposition 8.15] by computing
$\bm{\mathsf{A}}_{ji}=\frac{\partial}{\partial q_{j}}P_{i}$ (17)
The vectors $\{\bm{\mathsf{A}}_{j1},\ldots,\bm{\mathsf{A}}_{j,N-M}\}$ are not pairwise orthogonal but are orthogonal to $\mathcal{M}$, and thus can form a basis of the normal space at $\bm{q}$. We can write the eigenvalue problem using an arbitrary vector $b_{i}$ as
@@
This is a system of $N$ equations with $N-M$ unknowns. We can eliminate redundant equations by contracting each side with $\bm{\mathsf{A}}_{ij}$ to finally get
$\lambda(\bm{\mathsf{A}}^{T}\bm{\mathsf{A}})\bm{b}=(\bm{\mathsf{A}}^{T}\bm{\mathsf{B}}\bm{\mathsf{A}})\bm{b}$ (19)
@@
IV Calculation for the trimer problem
Figure 2: (a) Parametrisation used for the trimer uses the position of the central particle ($b$), bond angle $\psi$ and bond lengths $l_{a}$ and $l_{b}$, with identical equilibrium spring lengths $l_{0}$. Rigid-body rotation around the particle $b$ is omitted for clarity. (b) Motion along the symmetric (top) and antisymmetric eigenvectors of the hessian $\bm{\mathsf{H}}$ of the harmonic confinement function. Note that the central bead must move for the overall motion to be orthogonal to the constraining manifold $\mathcal{M}$. (c) Equilibrium probability density function (PDF) for the bond angle $\psi$ from the Langevin numerical simulation. Solid line is the theoretical prediction of Eq. (36).
@@ IV Calculation for the trimer problem
where $N$ is a normalisation constant. Therefore, the marginal density of the bond angle is proportional to $\sin\psi$ as expected. However, this is simply not the case of a uniform distribution––that distribution corresponds to $|\bm{\mathsf{H}}|=\text{const}$. This can be examined in greater detail by looking at the eigenvectors of $\bm{\mathsf{H}}$, visualised in Fig. 2(b). There are two eigenvectors which correspond to symmetric and antisymmetric motion of the terminal beads. Counterintuitively, the vectors in $\mathcal{M}^{\perp}$ change the orientation of at least one bond when the bond length changes.
@@
We additionally corroborate this result with a Brownian Dynamics simulation in Python using the package pychastic [16]. Choosing $k_{B}T$ as the energy scale, $l_{0}$ as the distance scale and $l_{0}^{2}/D$, as the timescale, with $D$ being the diffusion coefficient, the corresponding stochastic BD equation takes the form
@@ V Cyclic tetramer
which is not non-normalizable near 0 and $\pi$. Physically, as we approach the stiff spring limit, the tetramer spends more and more time in the folded state, and unless some repulsive potential is added, the molecule in the stiff spring limit looks predominantly like a trimer, alternating between beads $(a,c)$ and $(b,d)$ coinciding in space.
@@ VI Conclusions
This shows that the reasoning presented by Frenkel & Smit [3] cannot be universally applied. First, the dynamics and the distribution of a polymer connected with (stiff or soft) springs is independent of the mass of the monomers in the overdamped regime, and thus the limiting distribution cannot depend on the mass of the monomers. Second, putting aside the question of existence of the rigid rod distribution (or the lack thereof), the supposition that the stiff spring limit should always be the same, regardless of the details of the confining potential is clearly not true, and harmonic potentials will lead to a different outcome than springs that realise uniform confinement around the constraining manifold, as previously remarked, e.g., by van Kampen & Lodder [12], although without a general mathematical formulation. The difference can be arbitrarily large, as shown in the tetramer example. Frenkel & Smit remark in their book that for the bond length constraints of the type most often used in Molecular Dynamics simulations, the effect of hard constraints on the distribution functions seems to be relatively small [3, Chapter 15.1.1]. We believe this statement to be misleading. In the discussed case of the trimer molecule, the effect might indeed be small, but we have shown that for the tetramer system, which is still a relatively simple setup, the effect is pronounced. To provide a practical insight, we propose a straightforward technique for the analysis of the limiting distributions once the confining potential is specified in detail.

### v2

@@ III.1 Calculation of the shape term
Since the eigenvectors of $\bm{\mathsf{B}}$ are orthogonal and zero eigenvectors lie inside the tangent space $T\mathcal{M}$, we can calculate $|\bm{\mathsf{H}}|$ from the product of the non-zero eigenvalues of $\bm{\mathsf{B}}$. We can use knowledge of the tangent space of $\mathcal{M}$ to find the eigenvalues of $\bm{\mathsf{B}}$. If the confining function is of the form $W(\bm{q})=\sum_{i}P_{i}^{2}$, we can find vectors that lie in the normal space $\mathcal{M}^{\perp}$ [17, proposition 8.15] by calculating
$\bm{\mathsf{A}}_{ji}=\frac{\partial}{\partial q_{j}}P_{i},$ (19)
The vectors $\{\bm{\mathsf{A}}_{j1},\ldots,\bm{\mathsf{A}}_{j,N-M}\}$ are not pairwise orthogonal, but they are orthogonal to $\mathcal{M}$, and thus can form a basis of the normal space at $\bm{q}$. We can write the eigenvalue problem using an arbitrary vector $\bm{b}$ as
@@
This is a system of $N$ equations with $N-M$ unknowns. We can eliminate redundant equations by contracting each side with $\bm{\mathsf{A}}_{lk}$ to finally get
$\lambda(\bm{\mathsf{A}}^{T}\bm{\mathsf{A}})\bm{b}=(\bm{\mathsf{A}}^{T}\bm{\mathsf{B}}\bm{\mathsf{A}})\bm{b},$ (21)
@@
IV Calculation for the trimer
@@
where $N$ is a normalisation constant. Therefore, the marginal density of the bond angle is proportional to $\sin\psi$ as expected. However, this is simply not the case of a uniform distribution––that distribution corresponds to $|\bm{\mathsf{H}}|=\text{const}$ and the notion of uniformity is inherited from the canonical measure in the canonical (Cartesian) coordinates. This can be examined in greater detail by looking at the eigenvectors of $\bm{\mathsf{H}}$, visualised in Fig. 2(b). There are two eigenvectors which correspond to symmetric and antisymmetric motion of the terminal beads. Counterintuitively, the vectors in $\mathcal{M}^{\perp}$ change the orientation of at least one bond when the bond length changes.
@@
We additionally corroborate this result with a BD simulation in Python using the package pychastic [18]. Choosing $k_{B}T$ as the energy scale, $l_{0}$ as the distance scale and $l_{0}^{2}/D$, as the timescale, with $D$ being the diffusion coefficient, the corresponding stochastic BD equation takes the form
@@ V Cyclic tetramer
which is non-normalizable near 0 and $\pi$. Physically, as we approach the stiff spring limit, the tetramer spends more and more time in the folded state, and unless some repulsive potential is added, the molecule in the stiff spring limit looks predominantly like a trimer, alternating between beads $(a,c)$ and $(b,d)$ coinciding in space.
@@ VI Conclusions
This shows that the reasoning presented by Frenkel & Smit [3] cannot be universally applied. First, the dynamics and the distribution of a polymer connected with (stiff or soft) springs is independent of the mass of the monomers in the overdamped regime, and thus the limiting distribution cannot depend on the mass of the monomers. Second, putting aside the question of existence of the rigid rod distribution (or the lack thereof), the supposition that the stiff spring limit should always be the same, regardless of the details of the confining potential is clearly not true, and harmonic potentials will lead to a different outcome than springs that realise uniform confinement around the constraining manifold, as previously remarked, e.g., by van Kampen & Lodder [13], although without a general mathematical formulation. The difference can be arbitrarily large, as shown in the tetramer example. Frenkel & Smit remark in their book that for the bond length constraints of the type most often used in MD simulations, the effect of hard constraints on the distribution functions seems to be relatively small [3, Chapter 15.1.1]. We believe this statement to be misleading. In the discussed case of the trimer molecule, the effect might indeed be small, but we have shown that for the tetramer system, which is still a relatively simple setup, the effect is pronounced. To provide a practical insight, we propose a straightforward technique for the analysis of the limiting distributions once the confining potential is specified in detail.